Grade 12th pass9 grade science

(a) Describe Thomson's model of the atom. Which subatomic particle was not present in Thomson's model of the atom? (b) The mass number of an element is 18. It contains 7 electrons. What is the number of protons and neutrons in it? What is the atomic number of the element?

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Jayesh Vashishtha
4 Years ago
Dear student,
 
(a) J.J Thomson proposed a model of atom that is similar to a water melon.
The electron in the sphere is considered as seeds in watermelon. 
He proposed that:
1. Atom consists of a positively charged sphere and the electrons are embedded in it.
2. The negative and positive charges are equal in magnitude so atom is electrically neutral.
This model did not tell us about the presence of neutrons in the atom.
 
(b) When mass no. is 18 and no. of electrons is 7 then
(i) No. of protons = 7
(ii) No. of neutrons = 18 - 7 = 11
(iii) Atomic no. = 7
 
 
reagards,
team Askiitians